In Fig. 30-63, R = 2.0 ohm and the ideal battery h
Answer #1

I = I0*(1-e-tR/L )

at t =5sec

I0 = E/R = 10/2 = 5 amp

1 = 5*(1-e^(-2*5/L)

=> -0.8 = -e^(-2*5/L)

=> 0.223 = 10/L

=> L = 10/(0.223) = 44.84 H